# Towards a Physics Foundation Model

### v2

@@ Abstract
Foundation models have revolutionized natural language processing through a “train once, deploy anywhere” paradigm, where a single pre-trained model adapts to countless downstream tasks without retraining. Access to a Physics Foundation Model (PFM) would be transformative—democratizing access to high-fidelity simulations, accelerating scientific discovery, and eliminating the need for specialized solver development. Yet current physics-aware machine learning approaches remain fundamentally limited to single, narrow domains and require retraining for each new system. We present the General Physics Transformer (GPhyT), trained on 1.8 TB of diverse simulation data, that demonstrates foundation model capabilities are achievable for physics. Our key insight is that transformers can learn to infer governing dynamics from context, enabling a single model to simulate fluid-solid interactions, shock waves, thermal convection, and multi-phase dynamics without being told the underlying equations. GPhyT achieves three critical breakthroughs: (1) superior performance across multiple physics domains, outperforming specialized architectures by up to 29x, (2) zero-shot generalization to entirely unseen physical systems through in-context learning, and (3) stable long-term predictions through 50-timestep rollouts. By establishing that a single model can learn generalizable physical principles from data alone, this work opens the path toward a universal PFM that could transform computational science and engineering.
@@ 1 Introduction
We envision a similar paradigm for physics-aware machine learning (PAML), where a single foundation model can simulate a wide range of physical systems, boundary conditions, and initial states. This allows end-users to employ the model for their individual use case without the need for extensive retraining or fine-tuning. However, current state-of-the-art physics models, such as physics-informed neural networks (PINNs) (Raissi et al., 2019), neural operators (Kovachki et al., 2023; Lu et al., 2021), and physics-aware recurrent convolutions (PARC) (Nguyen et al., 2023a; 2024), are fundamentally limited to solving a single, narrowly scoped physical system: While they excel at their specific task, they cannot generalize to new physics or boundary conditions without new or additional training. Even recent multi-physics approaches largely rely on fine-tuning or meta-learning, which still demands new data and training for each application (Penwarden et al., 2023; Cho et al., 2023; McCabe et al., 2024).
@@
Here, we propose that the path towards PFMs lies in emulating the in-context learning abilities of LLMs (Brown et al., 2020; Agarwal et al., 2024). Instead of being explicitly told the governing equations, a model should infer the underlying dynamics from a "prompt" consisting of a short sequence of prior states. Such a model could adapt its predictions on the fly, enabling a single, unified architecture to tackle a wide array of physical scenarios. This paradigm shift presents fundamental research challenges that we address through the General Physics Transformer (GPhyT) trained on a diverse 1.8 TB corpus of simulation data. With this model, we investigate three critical questions:
@@ 1. Q1:
    Can a single, large-scale transformer effectively model a wide range of disparate physical systems (e.g., incompressible flow, shock waves, convection) without any explicit, physics-describing features? (Section 4.1)
@@ 2 Related work
In physics, the pursuit of foundation models has largely focused on enhancing the generalization of neural surrogates. Researchers have explored meta-learning (Penwarden et al., 2023) and transfer learning (Subramanian et al., 2023; Goswami et al., 2022) to adapt pretrained models to new PDE systems with fewer data samples. Recently, multi-tasks models trained on multiple physical systems were explored McCabe et al. (2024). However, while these groups demonstrated superior accuracy compared to single-physics models, all opted for finetuning to unseen tasks. These efforts represent important progress, but they still fall short of the “train once, deploy anywhere” paradigm that we envision.
@@ 3.1 Architecture
In this study, we choose the first-order Forward Euler method. An ablation study in Appendix 6.4 showed that higher-order integrators like Runge–Kutta 4 provided no significant accuracy benefit for our framework, making Forward Euler the most computationally efficient choice.
@@ 4.1 Multi-physics learning
Figure 2: Average and median mean squared error (MSE) across all test sets for each model (lower is better). Number of parameters (in million) are given b elow model names.
@@
To answer our first research question Q1—whether a single model can effectively learn to represent numerous, disparate physical systems—we evaluated GPhyT’s single-step prediction accuracy across our entire multi-physics test set. We benchmarked against two established architectures: the Fourier Neural Operator (FNO) (Kovachki et al., 2023), and a standard UNet, a widely-adopted baseline for spatiotemporal prediction tasks. All models were trained identically on 4 initial frames to predict the subsequent frame, ensuring a fair comparison.
Figure 2 presents the average and median mean squared error (MSE) across all test sets. GPhyT demonstrates substantial performance gains in average MSE and achievs a 5× reduction in median MSE compared to UNet and a 29× reduction compared to FNO at equivalent model sizes (size M). Detailed per-dataset losses are provided in SI section 6.2. Remarkably, even our smallest variant (GPhyT-S) with only 10% of the parameters outperforms the full-sized FNO and UNet when comparing median MSE. This efficiency gain suggests that the spatiotemporal attention mechanism captures cross-scale physical interactions more effectively than traditional convolutional or neural operator approaches. Furthermore, GPhyT exhibits favorable scaling properties with performance gains from the S to XL variants, indicating potential for further improvements with increased model capacity.
Figure 3: Next timestep prediction (velocity magnitude) of GPhyT, Unet, and FNO in their M-variant against ground truth (GT). Top row: Incompressible flow with periodic boundary conditions and $\Delta t=8$. 2nd row: Shear flow ($\Delta t=4$). 3rd row: Euler shock waves with periodic boundary conditions ($\Delta t=4$). 4th row: Rayleigh–Bénard convection ($\Delta t=4$). The results can be best viewed on a high-definition digital monitor.
Figure 3 provides qualitative comparisons across representative physical systems, categorized by complexity. For smooth systems (top two rows: obstacle flow and shear flow), GPhyT and UNet precisely capture fine-scale vortical structures and boundary layer dynamics, while the FNO struggles with spatial localization. For more challenging systems featuring discontinuities and chaotic dynamics (bottom two rows: Euler shocks and Rayleigh–Bénard convection), the performance gap widens dramatically. GPhyT maintains sharp shock fronts and accurately tracks convective plumes, while the FNO fails to resolve these critical features, producing overly smoothed predictions that miss essential physical phenomena. The UNet produces reasonable approximations but exhibits some diffusion of sharp features. These results suggest that the GPhyT architecture, with its ability to dynamically attend to relevant spatiotemporal features, is inherently better suited as a Physics Foundation Model than FNO and UNet.
@@ 4.2 In-context learning
Table 2: Average MSE for in-context learning tasks. The model is evaluated on systems with new boundary conditions and entirely new physics, with known systems provided as a baseline.
Dataset Variant Zero-shot MSE
Obstacle flow
   Symmetric Known $2.97\times 10^{-4}$
   Periodic Known $5.32\times 10^{-4}$
   Open New $8.01\times 10^{-3}$
Euler
   Periodic Known $8.92\times 10^{-3}$
   Open New $9.09\times 10^{-3}$
Turbulent layer New $2.05\times 10^{-1}$
Supersonic New $1.40\times 10^{-1}$
@@
The defining characteristic of a true foundation model is its ability to adapt to new tasks without additional training—a capability that fundamentally distinguishes them from traditional specialized models. In language models, this emerges through in-context learning, where models leverage prompts to perform tasks never explicitly seen during training (Brown et al., 2020). To investigate our research question Q2, whether GPhyT exhibits similar emergent capabilities for physics, we designed increasingly challenging generalization experiments: First, we evaluated the model on systems with modified boundary conditions completely absent from the training data. Second, we pushed the boundaries further by presenting entirely novel physical phenomena, including supersonic flows and turbulent radiative layers never encountered during training. These experiments probe whether the model has learned transferable physical principles rather than memorizing dataset-specific patterns.
Figure 4: Qualitative results for in-context learning. GPhyT accurately predicts the next state for an incompressible flow and Euler flow with unseen open boundary conditions (1st and 2nd column). For the completely new physics of supersonic flow (3rd column), it successfully captures the formation of a bow shock wave, demonstrating strong zero-shot generalization. For the turbulent radiative layer (4th column), general turbulent features are present while finegrained details are missing. The results can be best viewed on a high-definition digital monitor.
The quantitative results are summarized in Table 2, with qualitative examples shown in Figure 4 and Section 6.8 in the Appendix. For incompressible flow around obstacles, introducing a new "open" boundary condition, where mass is not conserved, results in an average MSE nearly identical to the MSE achieved on the "known" symmetric system. Even more striking is the result for the chaotic Euler system, where the open boundary variant achieves the same accuracy compared to the known periodic system. Thus, GPhyT successfully infers new boundary conditions from the prompt without the need for finetuning.
@@
The most challenging tests are two completely new physical systems: Supersonic flow around an obstacle and a turbulent radiative layer. Here, the MSE is significantly higher than in the known systems, emphasizing the significant challenge. Yet, Figure 4 shows that GPhyT still captures the essential dynamics, correctly forming a bow shock and the overall flow field for the supersonic flow. Similarly, the general turbulent features are present in the turbulent radiative layer. Naturally, finetuning the model on the new systems as done in previous studies McCabe et al. (2024); Subramanian et al. (2023) will yield even more accurate predictions, however, the ability to extrapolate and produce a physically plausible result for entirely new physics, even with reduced accuracy, is a powerful demonstration of emergent generalization.
4.3 Long-range Prediction
As stated in research question Q3, the true utility of any physics surrogate model is measured by its ability to maintain stability and accuracy over extended temporal horizons. This task is exceptionally challenging, as it requires the model to generate a full trajectory from an initial state, with prediction errors from each step accumulating over time. It is a critical test of a model’s physical consistency and a common failure point that is often omitted in literature.
Figure 5: (a) Autoregressive rollout prediction for velocity magnitude of GPhyT-XL on the Euler shockwaves with periodic boundary conditions. (b) Average mean squared error (MSE) over 50 time steps for all known physical systems. (c) Average MSE over 50 time steps for the novel unseen physical systems.
Figure 5a) illustrates an autoregressive rollout for an Euler shockwave with known periodic boundary conditions, exhibiting low error accumulation. In the initial phase (t=16), GPhyT-XL maintains high fidelity, accurately capturing the intricate vortex structures and shock fronts of the ground truth. As the rollout progresses, error accumulation becomes visible as a loss of high-frequency detail (t=32). However, the model successfully preserves the global dynamics and physical plausibility of the flow; the large-scale structures and boundary interactions remain stable and coherent. Similar trajectories for additional systems are presented in Appendix 6.9. The quantitative evaluation across the known physical systems is shown in Figure 5b). For most physics, GPhyT’s exhibit near-linear error accumulation. Exceptions are twophase flow, where changes can occur extremely abrupt, and shear flow, where a chaotic initial system converges to a quasi-static solution.
Extension to out-of-distribution systems (Figure 5c) provides further insight into the model’s generalization mechanisms. Systems with modified boundary conditions maintain error trajectories comparable to their training counterparts. Specifically, Euler with open boundaries tracks the periodic case almost identically, suggesting the model has disentangled bulk dynamics from boundary effects. Similarly, for entirely novel physics error growth rates are higher, highlighting the need for even more accurate and stable physics foundation models.

### v3

@@ Abstract
Foundation models have revolutionized natural language processing through a “train once, deploy anywhere” paradigm, where a single pre-trained model adapts to countless downstream tasks without retraining. Access to a Physics Foundation Model (PFM) would be transformative—democratizing access to high-fidelity simulations, accelerating scientific discovery, and eliminating the need for specialized solver development. Yet current physics-aware machine learning approaches remain fundamentally limited to single, narrow domains and require retraining for each new system. We present the General Physics Transformer (GPhyT), trained on 1.8 TB of diverse simulation data, that demonstrates foundation model capabilities are achievable for physics. Our key insight is that transformers can learn to infer governing dynamics from context, enabling a single model to simulate fluid-solid interactions, shock waves, thermal convection, and multi-phase dynamics without being told the underlying equations. GPhyT achieves three critical breakthroughs: (1) superior performance across multiple physics domains, outperforming specialized architectures by more than 7x, (2) plausible zero-shot generalization to entirely unseen physical systems through in-context learning, and (3) more stable long-term predictions through long-horizon rollouts. By establishing that a single model can learn generalizable physical principles from data alone, this work opens the path toward a universal PFM that could transform computational science and engineering.
@@ 1 Introduction
We envision a similar paradigm for physics-aware machine learning (PAML), where a single foundation model can simulate a wide range of physical systems, boundary conditions, and initial states. This allows end-users to employ the model for their individual use case without the need for extensive retraining or fine-tuning. However, current state-of-the-art physics models, such as physics-informed neural networks (PINNs) (Raissi et al., 2019), neural operators (Kovachki et al., 2023; Lu et al., 2021), and physics-aware recurrent convolutions (PARC) (Nguyen et al., 2023a; 2024), are fundamentally limited to solving a single, narrowly scoped physical system: While they excel at their specific task, they cannot generalize to new physics or boundary conditions without new or additional training. Even recent multi-physics approaches largely rely on fine-tuning or meta-learning, which still demands new data and training for each application (Penwarden et al., 2023; Cho et al., 2023; McCabe et al., 2024; Cao et al., 2025; Hao et al., 2024; Herde et al., 2024; Morel et al., 2025).
@@
Here, we propose that the path towards PFMs lies in emulating the in-context learning abilities of LLMs (Brown et al., 2020; Agarwal et al., 2024). Instead of being explicitly told the governing equations, a model should infer the underlying dynamics from a "prompt" consisting of a short sequence of prior states. Such a model could adapt its predictions on the fly, enabling a single, unified architecture to tackle a wide array of physical scenarios. Furthermore, the model should use little implicit bias to remain as general as possible. This paradigm shift presents fundamental research challenges that we address through the General Physics Transformer (GPhyT) trained on a diverse 1.8 TB corpus of simulation data. With this model, we investigate three critical questions:
@@ 1. Q1:
    Can a single, large-scale (but simple) transformer effectively model a wide range of disparate physical systems (e.g., incompressible flow, shock waves, convection)? (Section 4.1)
@@ 2 Related work
In physics, the pursuit of foundation models has largely focused on enhancing the generalization of neural surrogates. Researchers have explored meta-learning (Penwarden et al., 2023; Morel et al., 2025) and transfer learning (Subramanian et al., 2023; Goswami et al., 2022) to adapt pretrained models to new PDE systems with fewer data samples. Recently, multi-tasks models trained on multiple physical systems were explored. These models either use multiple (McCabe et al., 2024; Hao et al., 2024; Cao et al., 2025; Nguyen et al., 2025) or single time steps (Herde et al., 2024) as input and autoregressively predict the temporal evolution of the system. However, while these groups demonstrated superior accuracy compared to single-physics models, all opted for finetuning to unseen tasks. These efforts represent important progress, but they still fall short of the “train once, deploy anywhere” paradigm that we envision.
@@ 3.1 Architecture
In this study, we choose the first-order Forward Euler method, however more accurate integrators like Heun or Runge–Kutta 4 can be chosen as well.
@@ 4.1 Multi-physics learning
Figure 2: Median normalized mean square error (NMSE) of all models across the test datasets for next step prediction. Losses are grouped for each dataset and the overall loss. The error bars indicate the 25th and 75th percentile errors. GPhyT shows the overall lowest error and lowest in all but one dataset.
@@
To address our first research question Q1—whether a single general model can effectively learn to represent numerous, disparate physical systems—we evaluated GPhyT’s single-step prediction accuracy across our entire multi-physics test set. We benchmarked against multiple established baseline models designed for multi-physics predictions: MPP (McCabe et al., 2024), DPOT (Hao et al., 2024), and Poseidon (Herde et al., 2024). All three are transformer-based architectures that have demonstrated state-of-the-art performance on learning multiple physical systems simultaneously. Both MPP and DPOT use multiple input frames (similar to GPhyT), while Poseidon uses only a single input timestep. Additionally, we included a standard UNet as a baseline without implicit physics-specific biases. All models were trained identically to predict the subsequent frame, ensuring a fair comparison.
Figure 2 presents the median normalized mean squared error (NMSE) for all test datasets as well as the overall error. We report the median rather than the mean, as all models experience occasional severe mispredictions that heavily skew the mean error, likely caused by the normalization of the error and select few extremely challenging samples. Overall, GPhyT demonstrates substantial accuracy gains over all baselines, with the next-best model (DPOT) exhibiting $7\times$ higher NMSE. Notably, the UNet also achieves lower error than the more sophisticated multi-physics models Poseidon and MPP.
Examining the per-dataset results reveals that different physical systems present vastly different challenges. Incompressible, steadily moving fluids (incompressible Navier-Stokes and shear flow) yield the lowest errors across all models. More complex systems involving heat transfer, such as Rayleigh-Bénard convection, or the Euler dataset, which features shockwaves and sharp discontinuities, present greater challenges and result in higher errors. This is expected, as such phenomena are notoriously difficult to resolve accurately. Nevertheless, GPhyT demonstrates robust performance across all tasks, achieving the lowest NMSE on all physics except Euler. While the three leading models (GPhyT, UNet, and DPOT) show a steady increase in NMSE for more challenging tasks, the NMSE for Poseidon and MPP remains nearly constant – potentially indicating difficulties in adapting to varying systems. It is important to note that all models still exhibit high variability in their prediction accuracy, as illustrated by the large error bars. These results suggest that the GPhyT architecture, despite its generality and simplicity, is inherently better suited as a Physics Foundation Model than other multi-physics approaches. Its ability to dynamically attend to relevant spatiotemporal features appears to be a key advantage for handling diverse physical phenomena within a single model.
4.2 Long-range Prediction
As stated in research question Q2, the true utility of any physics surrogate model is measured by its ability to maintain stability and accuracy over extended temporal horizons. This task is exceptionally challenging and remains unsolved, as it requires the model to generate a full trajectory from an initial state, with prediction errors from each step accumulating over time. Long-range prediction constitutes a critical test of a model’s physical consistency and represents a common failure point that is often omitted in the literature.
Figure 3: (a) Overall autoregressive long-horizon prediction (median NMSE) for all models on the known physical systems up to 24 prediction steps. (b) Visualization of prediction step t=1 and t=24 of Euler shockwaves with ground truth (GT), the worst model (MPP) and the two best (Poseidon, GPhyT). The images can be best viewed on a high-definition digital monitor.
Figure 3a shows the median NMSE as a function of autoregressive prediction steps. During the initial phase of the rollout (t=1t=188), all models exhibit significant error accumulation, which continues at a similar rate for longer time horizons. Interestingly, Poseidon shows less initial error accumulation but starts from already elevated error levels. Compared to the baselines, GPhyT performs exceptionally well, only surpassed slightly by Poseidon in the later stages of the rollout. Inspection of 12-step and 24-step rollout NMSE for individual datasets (see Appendix 6.1) reveals that Poseidon and GPhyT are competitive across most datasets, with the exception of shear flow, where Poseidon achieves substantially lower error rates, albeit with significant error bars. Since the slopes are similar across all models, we hypothesize that the initial error is of critical importance. Furthermore, error accumulation appears to be an inherent property of all neural architectures, suggesting that hybrid models or alternative approaches may be necessary to mitigate this limitation.
To investigate the failure modes of the models, we visualized representative prediction samples from the best-performing models as well as the worst (MPP) in Figure 3b and Appendix 6.4. Upon close inspection, Poseidon is unable to reproduce high-frequency details such as Euler shockwaves, whereas GPhyT successfully preserves the global dynamics and physical plausibility of the flow. MPP and DPOT, which perform worst in autoregressive rollouts, often fail to match the ground truth data entirely. Overall, it must be emphasized that no model is capable of high-fidelity predictions, let alone replacing numerical solvers. However, GPhyT shows not only the highest accuracy but also the most physical plausible and consistent predictions. Achieving the precision required for practical engineering applications will necessitate orders of magnitude improvement in both accuracy and stability.
@@ 4.3 In-context learning
The defining characteristic of a true foundation model is its ability to adapt to new tasks without additional training – a capability that fundamentally distinguishes foundation models from traditional specialized approaches. In language models, this emerges through in-context learning, where models leverage prompts to perform tasks never explicitly seen during training  (Brown et al., 2020). To investigate our research question Q3, whether GPhyT exhibits similar emergent capabilities for physics, we designed increasingly challenging generalization experiments. First, we evaluated the model on systems with modified boundary conditions that were completely absent from the training data. Second, we pushed the boundaries further by presenting entirely novel physical phenomena, including supersonic flows and turbulent radiative layers never encountered during training. These experiments probe whether the model has learned transferable physical principles rather than merely memorizing dataset-specific patterns.
Figure 4: (a) Overall autoregressive long-horizon prediction (median NMSE) for all models on the novel physical systems up to 24 prediction steps. (b) Separate graphs for autoregressive long-horizon prediction of systems with new boundary condations (top) and completely new physics (bottom).
The quantitative results are summarized in Figure 4. Additionally, predictions from all models are visualized in Appendix 6.5. Overall, GPhyT outperforms all other foundation models on the crucial long-horizon predictions. Notably, it is the only model that remains below an NMSE of 1, which is the threshold beyond which predicting the simple mean of the image would yield higher accuracy. The visualizations in the Appendix show that GPhyT and UNet, as the more general models, maintain greater stability and avoid diverging predictions. Figure 4b presents the same data split into systems with novel boundary conditions (top) and completely novel physical systems (bottom). For incompressible flow around obstacles and the complex Euler shockwaves with new boundary conditions, our model significantly outperforms the baselines. This demonstrates that GPhyT successfully infers new boundary conditions from the prompt alone, without requiring fine-tuning.
@@
The most challenging tests involve two completely new physical systems: supersonic flow around an obstacle and a turbulent radiative layer. Here, the initial NMSE is considerably higher than for the known systems, underscoring the significant challenge posed by truly out-of-distribution physics. Nevertheless, Figure 4b shows that GPhyT still achieves lower errors than all baselines. Fine-tuning the model on these new systems, as done in previous studies, would naturally yield even more accurate predictions (McCabe et al., 2024; Herde et al., 2024; Hao et al., 2024). However, the ability to extrapolate and produce physically plausible results for entirely new physics—even with reduced accuracy—constitutes a powerful demonstration of emergent generalization. This capability represents a fundamental step toward the “train once, deploy anywhere” paradigm that defines true foundation models.
4.4 Ablation: What Makes a Good Physics Foundation Model
To understand the key design choices that enable GPhyT’s strong performance, we conduct ablation studies examining architectural components, model scale, and input context length. Two design choices distinguish GPhyT from a standard video transformer: predicting the time derivative $\frac{\partial X}{\partial t}$ rather than the next state directly, and providing explicit spatial and temporal derivatives as additional input features. Figure 5a compares our neural differentiator framework against direct next-state prediction, showing that decoupling the learning of dynamics from numerical time integration yields substantially lower errors across all rollout horizons. We hypothesize that predicting derivatives provides a more natural learning target that generalizes better across different temporal scales. Figure 5b demonstrates the impact of explicit derivative features: without the computed spatial ($\partial_{x}$, $\partial_{y}$) and temporal ($\partial_{t}$) derivatives concatenated to the input, prediction accuracy degrades by nearly an order of magnitude for extended rollouts, as these features provide crucial local gradient information for resolving sharp gradients and discontinuities. Together, these results show that a general-purpose video transformer, augmented with lightweight physics-motivated modifications, can outperform specialized multi-physics architectures.
Figure 5: Ablation studies on the known datasets with median NMSE: (a) Comparing GPhyT against a version predicting the next state directly. (b) Comparing GPhyT against a version without explicit spatial and temporal derivatives as input. (c) Scaling behavior of GPhyT. (d) Effect of the number of input time steps on rollout performance.
Beyond architectural choices, GPhyT exhibits favorable scaling behavior (Figure 5c): increasing model capacity from 9M to 385M parameters yields consistent accuracy improvements, suggesting that further gains may be achievable with larger models. Finally, the number of input timesteps ($N_{\text{input}}$) directly controls the model’s capacity for in-context learning. Figure 5d shows that a single input timestep, which provides no temporal context, produces the highest errors, while the largest performance gain occurs when increasing from one to two timesteps. Additional timesteps continue to improve accuracy with diminishing returns, following a log-linear trend. This analysis has practical implications: while longer prompts improve accuracy, they increase computational cost quadratically due to self-attention and must be generated by a numerical solver at inference time. Our results suggest that $N_{\text{input}}=4$ provides a favorable accuracy-efficiency trade-off, requiring substantially fewer input frames than prior work (McCabe et al., 2024; Hao et al., 2024) while maintaining competitive performance.
4.5 Limitations
While GPhyT demonstrates promising advances toward true physics foundation models, several key limitations remain:
2D data constraints: Due to data scarcity and computational limitations, the current model, as well as most other multi-physics models are restricted to 2D systems. However, the proposed architecture is directly extensible to 3D systems, and the increased computational demands can be mitigated by employing larger temporal patch sizes.
Long-term stability: Although GPhyT achieves remarkable accuracy in long-term rollout predictions, it falls considerably short of the precision exhibited by numerical solvers. Significantly lower prediction errors are essential for practical engineering applications.
Limited physics coverage: GPhyT is currently trained exclusively on fluid dynamics and heat transfer systems. A comprehensive physics foundation model would require incorporation of diverse physical domains, including mechanics, chemistry, molecular dynamics, and optics.
Fixed domain resolution: The model is trained on 256×128 resolution images. While this resolution is adequate for many simulation scenarios, widespread adoption may necessitate a model capable of handling variable domain sizes and resolutions. However, per-dataset normalization allows GPhyT to train on multiple resolutions of the same system, effectively learning discretization-invariance.